# MemeIntel: Explainable Detection of Propagandistic and Hateful Memes

### v1

@@ Abstract
The proliferation of multimodal content on social media presents significant challenges in understanding and moderating complex, context-dependent issues such as misinformation, hate speech, and propaganda. While efforts have been made to develop resources and propose new methods for automatic detection, limited attention has been given to label detection and the generation of explanation-based rationales for predicted labels. To address this challenge, we introduce MemeXplain, an explanation-enhanced dataset for propaganda memes in Arabic and hateful memes in English, making it the first large-scale resource for these tasks. To solve these tasks, we propose a multi-stage optimization approach and train Vision-Language Models (VLMs). Our results demonstrate that this approach significantly improves performance over the base model for both label detection and explanation generation, outperforming the current state-of-the-art with an absolute improvement of $\sim 3$% on ArMeme and $\sim 7$% on Hateful Memes. For reproducibility and future research, we aim to make the MemeXplain dataset and experimental resources publicly available. [^1] Please contact the corresponding author.
@@ 1 Introduction
Despite the rapid growth of multimodal content—integrating images, text, and sometimes video—the automated detection of harmful and false information on online news and social media platforms has become increasingly critical. In particular, identifying propaganda and hate in memes is essential for combating misinformation and minimizing online harm. While most research has focused on textual analysis, multimodal approaches have received comparatively less attention. In propaganda detection, text-based methods have evolved from monolingual to multilingual setups Piskorski et al. (2023); Hasanain et al. (2023), initially through binary classification and later via multilabel and fine-grained span-level tasks Barrón-Cedeno et al. (2019); Habernal et al. (2017, 2018); Da San Martino et al. (2019). Hate speech detection has similarly progressed from text-based to multimodal approaches that integrate both textual and visual elements. Recent methods have shifted from transformer-based text detection Fortuna and Nunes (2018) toward techniques that incorporate visual context Kiela et al. (2020) by leveraging fusion strategies, attention mechanisms, and contrastive learning to boost accuracy, especially when hateful intent is conveyed through text-image interplay Alam et al. (2022).
@@
The emergence of LLMs has demonstrated significant capabilities across various disciplines. Consequently, efforts have been made to leverage Vision-Language Models (VLMs) Zhang et al. (2024) and prompting techniques to enhance the detection and classification of harmful and propagandistic memes Cao et al. (2023). LLM-based models utilize prompt-based learning Cao et al. (2022), contrastive learning techniques such as CLIP Kumar and Nandakumar (2022), and cross-modal attention mechanisms to better capture implicit hate and propaganda.
Despite significant progress, challenges remain in detecting implicit hatefulness, particularly when sarcasm or an ironic dissonance exists between text and images. Propagandistic memes further complicate detection by employing emotional appeals, humor, cultural references, manipulative language, and other rhetorical strategies. To address these nuances, it is crucial for a system to provide not only accurate predictions but also interpretable explanations that reveal the underlying reasoning behind its decisions Hee et al. (2023); Yang et al. (2023); Huang et al. (2023); Sun et al. (2023). Such explanations enhance classifier reliability by helping end users understand the decision, provided they maintain a natural tone and relate closely to the visual elements of the meme.
An explanation-based approach offers numerous advantages and enhances performance across various tasks Li et al. (2022); Magister et al. (2022); Nandi et al. (2024); Kumari et al. (2024). While most studies have focused on textual content Li et al. (2022); Magister et al. (2022), a few recent approaches Nandi et al. (2024); Kumari et al. (2024) have applied explainability to images. However, these methods rely on QA-based explanations that lack naturalness, use multiple inference calls with custom models—thereby increasing computational complexity—and employ explanations only during training rather than as an inference output. These limitations motivated us to explore a simplified procedure for meme classification and explanation generation. To overcome the above limitations, we propose a novel procedure that achieves state-of-the-art performance on the target tasks across two distinct datasets. Our contributions are briefly outlined below:
  • We developed explanation-enhanced datasets, MemeXplain, using a rapid and low-cost annotation procedure;
@@
  • We proposed an efficient multi-stage optimization procedure that significantly improves performance;
  • With the experiments we achieved state-of-the-art performance on two types of datasets related to propaganda and hateful content detection.
Our findings are as follows: (a) A higher human evaluation score suggests that explanations from stronger models (e.g., GPT-4o) are reliable and can serve as gold-standard explanations for training smaller models. (b) Task-specific fine-tuning improves performance over the base model. (c) Our multi-stage optimization approach benefits both label detection and explanation generation. Overall, our work is the first to enhance VLMs for simultaneous propaganda and hateful content detection while providing natural reasoning to end users.
@@ 2 Related Work
2.1 Multimodal Propagandistic Content
@@
Following the previous research for propaganda detection using textual content Da San Martino et al. (2019), Dimitrov et al. (2021b) introduced SemEval-2021 Task 6 focusing on persuasion techniques detection in both textual and visual memes. Subsequently, the focus has extended to the detection of multilingual and multimodal propagandistic memes Dimitrov et al. (2024). Glenski et al. (2019) studied multimodal disinformation content on social media platforms in multilingual settings. Similar multimodal work on Arabic involves the development of datasets and shared task for propaganda detection Alam et al. (2024b); Hasanain et al. (2024). For the detection problem, typical approaches include a fusion of textual and visual embedding and a classification head on top them Hasanain et al. (2024); Shah et al. (2024), graph attention network based approach for multimodal visual-textual objects Chen et al. (2024).
2.2 Multimodal Hate speech
@@
Similarly, there has been growing interest in detecting multimodal hate speech Kiela et al. (2020); Velioglu and Rose (2020); Hee et al. (2022). Due to the lack of resources, Kiela et al. (2020) developed a large-scale dataset for multimodal hate identification. This study advanced research in this area and emphasized the importance of integrating textual and visual features for effective detection. The issue has also been explored through a multi-task learning framework for identifying hate speech in memes using multimodal features Sharma et al. (2022). To further progress in this field, efforts have been made to develop resources for multiple languages, including Arabic Alam et al. (2024a), Bangla Hossain et al. (2022), and English Hee et al. (2023). A more detailed summary of these earlier efforts can be found in Sharma et al. (2022), which also highlights key challenges and outlines future research directions.
2.3 Training with Explanations
@@
Integrating reasoning or explainability capabilities to enhance LLM/VLM performance has been shown to be highly beneficial for various tasks across multiple domains Plaat et al. (2024). This approach has also proven effective for knowledge distillation and model compression Li et al. (2022); Magister et al. (2022), where explanations generated by large LLMs improve the performance and capabilities of smaller LLMs. In the context of hateful speech, toxicity detection, and sentiment analysis, it has led to significant advancements Yang et al. (2023); Huang et al. (2023); Sun et al. (2023). For example, in the hateful speech detection task, Hare Yang et al. (2023) employs Chain-of-Thought (CoT) reasoning, while Huang et al. (2023) utilizes Chain of Explanation (CoE). Their aim is to improve the effectiveness of LLM-based sentiment classifiers by leveraging reasoning capabilities. Likewise, Sun et al. (2023) introduced a technique called Clue and Reasoning Prompting (CARP), which incorporates both reasoning and keywords as clues to support the reasoning process. In the following subsections, we specifically examine approaches closely related to our task, focusing on those that have applied VLM-based methods for analyzing hateful or propagandistic memes.
@@
CoT is a widely recognized prompting technique that generates a chain of reasoning to derive answers. A recent comprehensive CoT-based meme analysis study is presented in Kumari et al. (2024), which proposed a framework based on text- and image-based entity-object relationships using a scene graph. They applied a hierarchical three-step CoT-based prompting strategy to guide the LLM in identifying Emotion, Target, and Context, using these elements to build a model for meme analysis. Another recent work, called SAFE-MEME Nandi et al. (2024), proposed two multimodal datasets and introduced a structured reasoning framework for hate speech detection in memes. They developed a CoT prompting-based framework that incorporates Q&A-style reasoning and hierarchical categorization to classify memes as hateful (explicit or implicit) or benign. However, they did not evaluate their approach using the popular Hateful Memes dataset, preventing a direct numerical comparison with their results.
One drawback of these CoT-based approaches is that they rely on multi-step reasoning, requiring multiple inferences with VLMs. Our approach differs from these CoT-based methods in the following ways: (a) we do not employ a complex multistep CoT approach, eliminating the need for multiple LLM inferences, which significantly improves computational efficiency and reduces costs. (b) we focus on providing explanations alongside classification, helping the end-users better understand the reasoning behind classification decisions, thereby increasing reliability.
@@
Hee et al. (2023) constructed a dataset providing explanations for hateful memes. However, unlike us, they focused solely on evaluating explanation generation and did not perform classification tasks. Despite the availability of their data, we do not use it due to the lack of naturalness. In particular, their explanations do not fully account for image content or image-centric contextual perspectives.
@@ 3.1 ArMeme
The ArMeme dataset aimed to address the scarcity of Arabic-language datasets for multimodal propaganda detection. It comprises approximately $\sim$6k Arabic memes collected from various social media platforms, each manually annotated to identify propagandistic content Alam et al. (2024b). This dataset has been collected from different social media platforms, filtered, cleaned and manually annotated with four labels such as Not propaganda, Propaganda, Not-meme and Other. Table 1 provides the distribution of the data splits. The memes with “Not propaganda” category covers over half of the dataset ($\sim$66%), followed by “Propaganda” and the distribution of “Not-meme“ and “Other“ classes are significantly smaller. This distribution highlights a substantial class imbalance, particularly between “Not propaganda” and the other categories.
@@ 3.2 Hateful Meme
The Hateful Memes dataset Kiela et al. (2020), is a benchmark designed to evaluate multimodal hate speech detection. It consists of $\sim$12k memes, combining both text and images, carefully curated to ensure that effective classification requires an understanding of both modalities. The dataset was created using a mix of synthetically generated memes and real-world examples, sourced from social media, while ensuring a balanced distribution of hateful and non-hateful content. A key feature of this dataset is the inclusion of benign confounders, where individual elements of a hateful meme—either the image or the text—are altered to make it non-hateful. This approach prevents unimodal models (which rely only on text or images) from achieving high performance, reinforcing the need for true multimodal understanding. In Table 2, we report the distribution of hateful meme dataset used for this study. Note that hateful meme dataset consists of two other splits (dev-seen and test-seen), here, we used unseen versions.
Class Label Train Dev-seen Test-seen Total
Not Hateful 5,481 253 510 6,244
Hateful 3,019 247 490 3,756
Total 8,500 500 1000 10,000
Table 2: Distribution of hateful meme dataset.
@@ 4 MemeXplain: Explanation Generation
In this research, we generate explanations for two different stages: (a) during existing dataset enhancement, which leverages an expert VLM (such as GPT) to generate high-quality explanations and (b) during training/inference with a smaller VLM (such as Llama-3.2 11b). Figure 1 illustrates these different stages. Mathematically, these two stages can be described by the functions $f(i,l)=e$ and $g(i)=(l,e)$, where $e$ denotes the explanation, $l$ is the label, and $i$ is the input image or meme. Specifically, $f(i,l)$ returns an explanation $e$ given both $i$ and $l$, whereas $g(i)$ generates both the label $l$ and the explanation $e$ from only the input $i$.
@@ 4.1 VLMs for Explanation Generation
Determining the optimal length for explanations is important for balancing informativeness and cognitive load Herm (2023). Shen et al. (2022) explored the relationship between explanation length and human understanding, finding that the shortest rationales are often ineffective. Recently, Wang et al. (2023) also studied the effect of explanation size and found that human evaluators are reluctant to read longer explanations. To achieve an optimal balance, we iteratively tested various explanation lengths and ultimately set a limit of 100 words.
@@
To utilize GPT-4o OpenAI (2023), we accessed the OpenAI API via Azure services. Though recently released o1 models have shown promising directions for complex reasoning, they were not accessible to us. For explanation generation, we employed zero-shot learning. To ensure reproducibility, we set the temperature value to zero.
4.2 Human Evaluation
@@
For manual annotation, we first prepared an annotation guideline for the annotators. Additionally, we developed annotation guidelines and a platform (see Appendix B and A, respectively).
Evaluation Setting. For the Arabic meme task, we recruited annotators who are native Arabic speakers and fluent in English, all holding at least a bachelor’s degree. Because of their fluency, they also handled the hateful meme task. We provided necessary training and consultation, and all had prior experience with similar tasks.
A total of six annotators participated in the evaluation. In line with institutional requirements, each signed a Non-Disclosure Agreement (NDA), and a third-party company managed their compensation at standard hourly rates based on location.
Quality Assessment
In Table 4, we summarize the quality assessment of the explanations. We used 5-point Likert scale for various human evaluation metrics, including informativeness, clarity, plausibility, and faithfulness. We compute the average of the Likert scale value for all evaluation metrics. We manually evaluated 359 and 202 random samples for ArMeme Arabic and English explanations while 200 random examples were evaluated for the Hateful meme dataset. The average agreement scores for the ArMeme dataset with Arabic explanations are 4.23, 4.38, 4.24, and 4.16 for faithfulness, clarity, plausibility, and informativeness, respectively, indicating high agreement across all evaluation metrics. However, for the English explanations of ArMeme, the faithfulness and plausibility scores are relatively. To better understand this issue, we plan to conduct further evaluations on another set of explanations. For the Hateful Memes dataset, the average Likert scale agreement scores range from 4.562 to 4.682.
Dataset Faithfulness Clarity Plausibility Informative
ArMeme (Ar) 4.23 4.38 4.24 4.16
ArMeme (En) 3.91 4.50 3.81 4.13
Hateful meme 4.56 4.65 4.63 4.68
Table 4: Average Likert scale value for each human evaluation metric across different sets of explanations.

### v2

@@ Abstract
The proliferation of multimodal content on social media presents significant challenges in understanding and moderating complex, context-dependent issues such as misinformation, hate speech, and propaganda. While efforts have been made to develop resources and propose new methods for automatic detection, limited attention has been given to jointly modeling label detection and the generation of explanation-based rationales, which often leads to degraded classification performance when trained simultaneously. To address this challenge, we introduce MemeXplain, an explanation-enhanced dataset for propagandistic memes in Arabic and hateful memes in English, making it the first large-scale resource for these tasks. To solve these tasks, we propose a multi-stage optimization approach and train Vision-Language Models (VLMs). Our results show that this strategy significantly improves both label detection and explanation generation quality over the base model, outperforming the current state-of-the-art with an absolute improvement of $\sim{1.4}$% (Acc) on ArMeme and $\sim{2.2}$% (Acc) on Hateful Memes. For reproducibility and future research, we aim to make the MemeXplain dataset and scripts publicly available. [^1] https://github.com/MohamedBayan/MemeIntel
@@ 1 Introduction
Despite the rapid growth of multimodal content—integrating images, text, and sometimes video—the automated detection of harmful and false information on online news and social media platforms has become increasingly critical. In particular, identifying propaganda and hate in memes is essential for combating misinformation and minimizing online harm. While most research has focused on textual analysis, multimodal approaches have received comparatively less attention. In propaganda detection, text-based methods have evolved from monolingual to multilingual setups Piskorski et al. (2023); Hasanain et al. (2023), from binary to fine-grained span-level tasks Barrón-Cedeno et al. (2019); Habernal et al. (2017, 2018); Da San Martino et al. (2019). Hate speech detection has similarly progressed from text-based to multimodal approaches that integrate both textual and visual elements Kiela et al. (2020); Alam et al. (2022); Sharma et al. (2022).
@@
The emergence of LLMs has demonstrated significant capabilities across various disciplines. To capture implicit hate and propaganda, LLMs utilize different techniques Cao et al. (2022); Kumar and Nandakumar (2022), such as prompt-based methods, contrastive learning, and cross-modal attention. Consequently, efforts have been made to leverage VLMs Zhang et al. (2024) to classify the harmful and propagandistic memes Cao et al. (2023). Despite progress, detecting implicit hate, especially with sarcasm or irony, remains challenging. Propagandistic memes add complexity through emotional appeals, humor, cultural cues, and manipulative language. These challenges are especially pronounced for English-centric multilingual models, which often struggle to capture nuances in non-English content due to cultural differences and linguistic variation. Specific challenges for Arabic often include cultural and political references that do not translate well into Western contexts, which makes it difficult for general-purpose models to detect accurately. For instance, memes about local conflicts or those using religious symbols often rely on subtle cues that are deeply rooted in cultural context. Arabic-focused datasets have made it clear that understanding these nuances is essential for accurate detection and meaningful explanations Alam et al. (2024b). To address these nuances, it is crucial for a system to provide not only accurate predictions but also clear, relatable explanations that align with the meme’s visual context and aid user understanding Hee et al. (2023); Yang et al. (2023); Huang et al. (2023); Sun et al. (2023).
An explanation-based approach offers numerous advantages and improves performance in various tasks Li et al. (2022); Magister et al. (2022); Nandi et al. (2024); Kumari et al. (2024). Although most studies have focused on textual content Li et al. (2022); Magister et al. (2022), some recent approaches Nandi et al. (2024); Kumari et al. (2024) have applied explainability to images. However, these methods rely on QA-based explanations that lack naturalness, use multiple inference with custom models, thus increasing computational complexity, and employ explanations only during training rather than as an inference output. To address these limitations, we propose a simplified method that delivers strong performance in meme classification and explanation generation across two datasets.
We summarize our main contributions below:
  • We developed an explanation-enhanced datasets, MemeXplain, using a rapid and low-cost annotation procedure;
@@
  • We proposed an efficient multi-stage optimization procedure that mitigates gradient conflicts and avoids catastrophic forgetting by formulating the optimization problem from the perspective of domain adaptation and task‑incremental learning.
  • We achieved state-of-the-art performance on two types of datasets related to propaganda and hateful content detection.
Our findings are as follows: (a) A higher human evaluation score suggests that explanations from stronger models (e.g., GPT-4o) are reliable and can serve as gold-standard explanations for training smaller models; (b) Task-specific fine-tuning improves performance over the base model; and (c) Our multi-stage optimization improves label detection and explanation quality by reducing gradient conflicts and avoiding catastrophic forgetting, achieving state-of-the-art performance. Overall, our work is the first to enhance VLMs for simultaneous propaganda and hateful content detection while providing natural reasoning to end users.
@@ 2 Related Work
Multimodal Propagandistic Content.
@@
Following the previous research for propaganda detection using textual content Da San Martino et al. (2019), Dimitrov et al. (2021b) introduced SemEval-2021 Task 6 focusing on persuasion techniques detection in both textual and visual memes. Subsequently, the focus has extended to the detection of multilingual and multimodal propagandistic memes Dimitrov et al. (2024). Similar multimodal work on Arabic involves the development of datasets and shared task for propaganda detection Alam et al. (2024b); Hasanain et al. (2024). For the detection problem, typical approaches include a fusion of textual and visual embedding and a classification head on top them Hasanain et al. (2024); Shah et al. (2024), graph attention network based approach for multimodal objects Chen et al. (2024).
Multimodal Hate speech.
@@
Similarly, there has been growing interest in detecting multimodal hate speech Kiela et al. (2020); Velioglu and Rose (2020); Hee et al. (2022). Due to the lack of resources, Kiela et al. (2020) developed a large-scale dataset for multimodal hate identification. This study advanced research in this area and emphasized the importance of integrating textual and visual features for effective detection. To further progress in this field, efforts have been made to develop resources for multiple languages, including Arabic Alam et al. (2024a), Bangla Hossain et al. (2022), and English Hee et al. (2023). A more detailed summary can be found in Sharma et al. (2022), which also highlights key challenges and outlines future research directions.
Training with Explanations.
@@
Integrating reasoning or explainable capabilities to enhance LLM/VLM performance has been shown to be highly beneficial for various tasks across multiple domains Plaat et al. (2024). This approach has also proven effective for knowledge distillation and model compression Li et al. (2022); Magister et al. (2022), where explanations generated by large LLMs improve the performance and capabilities of smaller LLMs. For example, in the hate-speech detection task, Hare Yang et al. (2023) employs Chain-of-Thought (CoT) reasoning, while Huang et al. (2023) utilizes Chain of Explanation (CoE). Their aim is to improve the effectiveness of LLM-based sentiment classifiers by leveraging reasoning capabilities. Sun et al. (2023) proposed Clue and Reasoning Prompting (CARP), which uses keywords and reasoning to guide explanation.
CoT-Based Approaches.
@@
CoT is a widely recognized prompting technique that generates a chain of reasoning to derive answers. Kumari et al. (2024) proposed a CoT-based framework for meme analysis using scene graphs to capture text- and image-based entity-object relationships. Their three-step prompting strategy guides the LLM to identify Emotion, Target, and Context for effective meme interpretation. SAFE-MEME Nandi et al. (2024) introduced two multimodal datasets and a CoT-based reasoning framework for meme-based hate speech detection, using Q&A-style prompts and hierarchical labels. However, their method was not evaluated on the Hateful Memes dataset, limiting direct comparison.
Comparison with Prior Work.
@@
A key limitation of CoT-based methods is their reliance on multi-step reasoning, requiring multiple VLM inferences. In contrast, our approach: (a) avoids complex CoT steps, improving efficiency and reducing cost; and (b) generates explanations alongside classifications to enhance transparency and reliability. Hee et al. (2023) constructed a dataset providing explanations for hateful memes. However, unlike us, they focused solely on evaluating explanation generation and did not perform classification tasks. Despite the availability of their data, we do not use it due to the lack of naturalness. In particular, their explanations do not fully account for image content or image-centric contextual perspectives.
@@ 3.1 ArMeme
The ArMeme dataset comprises approximately $\sim$6k manually annotated Arabic memes collected from various social media platforms Alam et al. (2024b). This dataset has been manually annotated with four labels such as Not propaganda, Propaganda, Not-meme and Other. Table 1 provides the distribution of the data splits. The memes with “Not propaganda” category covers over half of the dataset ($\sim$66%), followed by “Propaganda” and the distribution of “Not-meme“ and “Other“ classes are significantly smaller. This distribution highlights a substantial class imbalance, particularly between “Not propaganda” and the other categories.
@@ 3.2 Hateful Meme
The Hateful Memes dataset Kiela et al. (2020), is a benchmark designed to evaluate multimodal hate speech detection. It consists of $\sim$12k memes, combining both text and images, carefully curated to ensure that effective classification requires an understanding of both modalities. The dataset was created using a mix of synthetically generated memes and real-world examples, sourced from social media, while ensuring a balanced distribution of hateful and non-hateful content. In Table 2, we report the distribution of hateful meme dataset used for this study. Note that hateful meme dataset consists of two other splits (dev-seen and test-seen), here, we used unseen versions.
Class Label Train Dev-unseen Test-unseen Total
Not Hateful 5,481 340 1,250 7,071
Hateful 3,019 200 750 3,969
Total 8,500 540 2,000 11,040
Table 2: Distribution of hateful-meme dataset.
@@ 4 MemeXplain: Explanation Generation
In this research, we generate explanations for two different stages: (a) during existing dataset enhancement, which leverages an expert VLM (such as GPT-4o) to generate high-quality explanations and (b) during training/inference with a smaller VLM (such as Llama-3.2 11b). Mathematically, these two stages can be described by the functions $f(i,l)=e$ and $g(i)=(l,e)$, where $e$ denotes the explanation, $l$ is the label, and $i$ is the input image or meme (can also be associated text). Specifically, $f(i,l)$ is defined as the function that generates an explanation $e$ for dataset enhancement, given an image $i$ and its corresponding label $l$ as input; and $g(i)$ is defined as the function that generates a label-explanation pair $(l,e)$ during training and inference, given only the image content $i$ as input. In the data preparation phase, we use a VLM to generate explanations conditioned on both the input and the label using the function $f$. Next, in the training phase, the model learns to jointly predict the label and explanation from the input alone using the function $g$.
@@ 4.1 VLMs for Explanation Generation
Determining the optimal length for the explanations is important to balance informativeness and cognitive load Herm (2023). Shen et al. (2022) explored the relationship between the length of the explanation and human understanding, finding that the shortest rationales are often ineffective. Recently, Wang et al. (2023) also studied the effect of explanation size and found that human evaluators are less willing to read longer explanations. To achieve an optimal balance, we iteratively tested various explanation lengths and ultimately set a limit of 100 words.
@@
To utilize GPT-4o OpenAI (2023), we accessed the OpenAI API via Azure services. Although recently released o1 models have shown promising directions for complex reasoning, they were not accessible to us. For explanation generation, we used zero-shot inference. To ensure reproducibility, we set the temperature value to zero.
4.2 Manual Annotation of Explanation
@@
Annotation Setup. For manual annotation, we first prepared an annotation guideline and a platform (see Appendix B and A, respectively) for the annotators. For the Arabic memes, we recruited annotators who are native Arabic speakers and fluent in English, all holding at least a bachelor’s degree. Because of their fluency, they also handled the hateful meme. We provided necessary training and consultation, and all had prior experience with similar tasks. A total of six annotators participated in the evaluation. In line with institutional requirements, each signed a Non-Disclosure Agreement (NDA), and a third-party company managed their compensation at standard hourly rates based on their location.
Annotation Agreement for Explanation. In Table 4, we summarize the annotation agreement scores of the explanations. We used 5-point Likert scale for various human evaluation metrics, including informativeness, clarity, plausibility, and faithfulness. We compute the average Likert scale value (from three annotators) for all evaluation metrics. We manually evaluated complete test sets for both ArMeme and Hateful meme except ArMeme with English explanation. For the later case we randomly selected 200 random examples. This decision was made to reduce the human annotation cost. The average agreement scores for the ArMeme dataset with Arabic explanations are >4.5>4.5 out of 5 indicating high agreement in all evaluation metrics. However, for the English explanations of ArMeme, the faithfulness and plausibility scores are relatively lower. To better understand this issue, we plan to conduct further annotation on the complete test set of explanations. For the Hateful Memes dataset, the average Likert scale agreement scores range from 4.04 to 4.10 out of 5.
In addition, we also computed the annotation agreement on ordinal scales by adopting the agreement index $r^{*}_{wg(j)}$ James et al. (1984), which compares observed variance in ratings to the maximum possible variance under complete disagreement. For each item, the agreement score is computed as: $r^{*}_{wg(j)}=1-\frac{S_{X}^{2}}{\sigma^{2}_{\text{mv}}}$ where $S_{X}^{2}$ is the observed variance across annotators and $\sigma^{2}_{\text{mv}}$ is the maximum variance possible given the scale (computed as $\sigma^{2}_{\text{mv}}=0.5(X_{U}^{2}+X_{L}^{2})-[0.5(X_{U}+X_{L})]^{2}$, with $X_{U}=5$ and $X_{L}=1$ for a 5-point scale). The average agreement scores for ArMeme and Hateful memes are above 0.83 and 0.92, respectively, for all metrics. These values indicate strong agreement O’Neill (2017).
Dataset Faithfulness Clarity Plausiblity Informative
ArMeme Ar expl. 4.64 4.69 4.69 4.74
ArMeme En expl. 3.91 4.50 3.81 4.13
Hateful meme 4.01 4.18 4.04 4.10
Table 4: Average Likert scale value for each human evaluation (annotation) metric across different sets of explanations.